You are given a population with standard deviation of 8.6. Determine the sample size


needed to estimate the mean of the population with error of 0.5 at 99 percent confidence.

1
Expert's answer
2022-01-27T11:46:26-0500

We are given that,

"E=0.5\\\sigma=8.6\\\alpha=0.01"

To find the required value of the sample size "n" needed to estimate the mean of the population with error of 0.5, we use the formula below.

"n=({Z_{\alpha\over2}\times \sigma\over E})^2" where "Z_{\alpha\over2}=Z_{0.001\over2}=Z_{0.005}=2.575"

Replacing for the values above in the formula, we have,

"n=({2.575\times8.6\over0.5})^2=({22.145\over0.5})^2=44.29^2=1961.6041\approx 1962"

Therefore, the sample size needed to estimate the mean of the population with error of 0.5 at 99 percent confidence is "n=1962."